OpenStudy (anonymous):

find dy/dx by implicit differentiation of: 2x^2 + x + xy = 1 and cos(x) + sqrt(y) = 5

OpenStudy (raden):

2x^2 + x + xy = 1 derive 2x^2, it is 4x then derive x, it is 1. derive xy, get y + xy' and for right side, derivative of 1 it is 0 now can you simplify this

OpenStudy (anonymous):

$$\cos x + y^{1/2} = 5 \\ \frac{d}{dx}[\cos x + y^{1/2}] = \frac{d}{dx}[5] \\ -\sin x + \frac{1}{2}y^{-1/2} \frac{dy}{dx} = 0 \\$$ Solve for \( \frac{dy}{dx} \).